### v1

# On homology spheres of the trivial local equivalence class

Abstract.
In the homology cobordism group $\Theta_{\mathbb{Z}}^{3}$, it is not known if there is a non-trivial linear dependence between Seifert fibered spheres. Based on involutive Heegaard Floer theory, Hendricks, Manolescu, and Zemke introduced the local equivalence group $\mathfrak{I}$ with the homomorphism $h:\Theta_{\mathbb{Z}}^{3}\rightarrow\mathfrak{I}$. And it is possible to find a non-trivial linear dependence between the image of Seifert fibered spheres in $\mathfrak{I}$ using the work of Dai and Stoffregen. Therefore, it is interesting to ask if such a linear dependence in $\mathfrak{I}$ originates from $\Theta_{\mathbb{Z}}^{3}$. In this paper, we prove that this is not the case in general. We show that infinitely many homology spheres linearly dependent in $\mathfrak{I}$ are linearly independent even in the rational homology cobordism group $\Theta_{\mathbb{Q}}^{3}$, by employing the $r_{s}$-invariants from the filtered instanton Floer homology developed by Nozaki, Sato, and Taniguchi. Moreover, these homology spheres include homology spheres of the trivial local equivalence class of $\operatorname{Pin(2)}$-equivariant Seiberg–Witten Floer stable homotopy type.
1. Introduction
The homology cobordism group $\Theta^{3}_{\mathbb{Z}}$ of homology spheres, introduced by González-Acuña [GA70], has played a central role in the development of low-dimensional topology over the last five decades [Man18, Hom23, Şav24]. The algebraic structure of $\Theta^{3}_{\mathbb{Z}}$ is linked to several topological problems such as the three fibers conjecture [FS87, Law88], the disproof of the triangulation conjecture for topological manifolds in high dimensions [Man16], and so on. While it is known that the homology cobordism group contains an infinite rank summand by [DHST23], the existence of a torsion element is one of the most notable open problems about the structure of $\Theta^{3}_{\mathbb{Z}}$.
Seifert fibered spheres have been core objects to study the structure of $\Theta^{3}_{\mathbb{Z}}$. In particular, the subgroup $\Theta^{3}_{SF}$ generated by Seifert fibered spheres itself contains an infinite rank subgroup in $\Theta^{3}_{\mathbb{Z}}$ [Fur90, FS90]. Moreover, the aforementioned infinite rank summand [DHST23] lies also in $\Theta_{SF}^{3}$. On the other hand, there are several infinite families of Seifert fibered spheres known to bound contractible manifolds or homology balls [AK79, CH81, FS81, Fic84, KŞ25], namely trivial in $\Theta_{\mathbb{Z}}^{3}$. However, it is neither known in $\Theta_{\mathbb{Z}}^{3}$ whether there exists a non-trivial Seifert fibered sphere of finite order, nor, more generally, whether there exists any non-trivial linear dependence among them. More precisely, we ask:
Question 1.1.
Are distinct Seifert fibered spheres always linearly independent in $\Theta_{\mathbb{Z}}^{3}$, provided they do not bound homology balls?
Hendricks, Manolescu, and Zemke [HMZ18] defined the notion of the local equivalence class of an involutive Heegaard Floer chain complex [HM17] of a homology sphere. Such classes form an abelian group $\mathfrak{I}$, called the local equivalence group, together with the homomorphism:
$h:\Theta_{\mathbb{Z}}^{3}\rightarrow\mathfrak{I}.$
Let $\mathfrak{I}_{SF}$ be the image of $\Theta_{SF}^{3}$ under $h$. While $\mathfrak{I}_{SF}$ has been particularly studied in [DM19, DS19] and is known to be isomorphic to $\mathbb{Z}^{\infty}$, it is possible to find many linear dependences in $\mathfrak{I}_{SF}$. Therefore, this fact and Question 1.1 motivate to study the kernel of the restriction of $h$ to $\Theta^{3}_{SF}$. In this paper, we detect a substantial structural difference between $\Theta^{3}_{SF}$ and $\mathfrak{I}_{SF}$, providing affirmative evidence for Question 1.1.
Theorem 1.2.
There are infinitely many homology spheres of the trivial local equivalence class that are linearly independent in the rational homology cobordism group.
We remark that we obtain linear independence not only in $\Theta_{\mathbb{Z}}^{3}$, but also in the rational homology cobordism group $\Theta_{\mathbb{Q}}^{3}$ of rational homology spheres. Together with the argument in Lemma 3.3, Furuta’s result [Fur90] is sufficient to prove linear independence in $\Theta^{3}_{\mathbb{Z}}$ of our homology spheres, which will be constructed later. However, this does not directly imply linear independence in $\Theta_{\mathbb{Q}}^{3}$, since the kernel of $\Theta_{\mathbb{Z}}^{3}\rightarrow\Theta_{\mathbb{Q}}^{3}$ is known to be non-trivial [FS84, AL18, Şav20, Sim21].
Dai and Manolescu [DM19] introduced a notion called monotone graded subroots of Seifert fibered spheres to simply determine their local equivalence classes. By analyzing monotone graded subroots systematically, Dai and Stoffregen [DS19] proved that:
$\mathfrak{I}_{SF}=\langle h(B(n))\rangle\cong\mathbb{Z}^{\infty},$
where $B(0)$ is the Poincaré homology sphere $\Sigma(2,3,5)$ and $B(n)$ is the Brieskorn sphere $\Sigma(2n+1,4n+1,4n+3)$ with $n>0$.
Therefore, potential candidates for a negative answer to Question 1.1 arise from linear dependence relations with respect to this explicit basis $\{h(B(n))\}_{n\geq 0}$ of $\mathfrak{I}_{SF}$. We construct homology spheres in Theorem 1.2 by relying on the aforementioned results [DM19, DS19] in terms of monotone graded subroots.
We consider the families of the Brieskorn spheres:
$Y_{1}(n)=\Sigma(4n+1,6n+2,12n+1)\quad\text{and}\quad Y_{2}(n)=\Sigma(4n-1,6n-2,12n-1).$
Their monotone graded subroots and hence their local equivalence classes were computed by Karakurt and the second author [KŞ22]. By using the fact that every monotone graded subroot is decomposed into simpler ones [DS19, Theorem 1.1], we first find each linear expression of the local equivalence class of $Y_{1}(n)$ and $Y_{2}(n)$ with respect to $h(B(n))$. From these linear relations, we obtain infinitely many homology spheres lying in $\operatorname{ker}(h)$ as follows:
$\big{\{}Y_{1}(n)\ \#\ -B(2n)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}\cup\big{\{}Y_{2}(n)\ \#\ -B(2n-1)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}.$
To obtain their linear independence in $\Theta^{3}_{\mathbb{Q}}$, we employ the $r_{s}$-invariants from the filtered instanton Floer homology, introduced by Nozaki, Sato, and Taniguchi [NST24]. They extended the previous work of Donaldson [Don02], Fintushel and Stern [FS85, FS90], and Furuta [Fur90] in instanton Floer theory.
Note that Theorem 1.2 indicates that the $d$-invariant [OS03a], the $\underline{d\mkern-2.0mu}\mkern 2.0mu$- and $\bar{d}$-invariants [HM17], and the $\phi_{n}$-invariants for all $n\geq 1$ [DHST23] must vanish for homology spheres above. In other words, letting $Y$ be any homology sphere generated by the above set, we see that:
$\underline{d}(Y)=d(Y)=\overline{d}(Y)=\phi_{n}(Y)=0.$
We observe that the $d$-invariant is not a complete rational homology cobordism invariant. Other homology spheres of infinite order in $\Theta^{3}_{\mathbb{Q}}$ with vanishing $d$-invariants can be easily deduced from [NST24]. We also note that the $\bar{\mu}$-invariant [Neu80, Sie80] vanishes by the result of [DS19].
On the other hand, Stoffregen [Sto20] defined an analogous group in Seiberg–Witten Floer theory together with the homomorphism:
$\Theta^{3}_{\mathbb{Z}}\to\mathfrak{LE},$
where the latter group denotes the local equivalence group of the $\operatorname{Pin}(2)$-equivariant Seiberg–Witten Floer stable homotopy type. Moreover, Dai, Sasahira, and Stoffregen [DSS23] recently proved that two Seifert fibered spheres have the same image in $\mathfrak{LE}$ if and only if they have the same image in $\mathfrak{I}$.
Motivated from this result, we provide another infinite family using the following family of the Brieskorn spheres:
$Y_{3}(n)=\Sigma(8n+1,12n+1,24n+5).$
Since its monotone graded subroot is known to be the same as the one of $B(4n)$ by [KŞ22], the resulting homology spheres are as follows:
$\{Y_{3}(n)\ \#\ -B(4n)\}_{n\geq 1}.$
These homology spheres also fall under Theorem 1.2. Moreover, they have the trivial local equivalence class in $\mathfrak{LE}$ as well. In particular, the $\kappa$-invariant [Man14] and the $\kappa o_{i}$-invariants for $i=0,\ldots,7$ [Lin15] from Seiberg–Witten Floer stable homotopy type, and the $\delta$-invariant [Frø10], the $\alpha$-, $\beta$-, and $\gamma$-invariants [Man16], and the $\underline{\delta}$- and $\overline{\delta}$-invariants [Sto17b] from Seiberg–Witten Floer homology must vanish for these homology spheres. Namely, letting $Y$ be any homology sphere generated by the above set, we also have the following:
$\alpha(Y)=\beta(Y)=\gamma(Y)=\underline{\delta}(Y)=\delta(Y)=\overline{\delta}(Y)=\kappa(Y)=\kappa o_{i}(Y)=0.$
To the authors’ best knowledge, it is not known whether there exists a Seifert fibered sphere for Theorem  1.2. The reason is that it is difficult to compute the $r_{0}$-invariant in general when the Fintushel–Stern $R$-invariant [FS85] is negative. Moreover, Issa and McCoy [IM20] proved that for a Seifert fibered sphere $Y$ with $d(Y)=0$, $R(Y)$ is always negative. Therefore, we conclude the introduction with the following question:
Question 1.3.
Is there a Seifert fibered sphere $Y$ with $d(Y)=0$ that does not bound a rational homology ball?
Conventions
[…]
Acknowledgments
The authors would like to thank Marco Golla, JungHwan Park, and Masaki Taniguchi for helpful conversations. The authors are also grateful to Hayato Imori and Imogen Montague for reading the draft and providing detailed feedback. The first author is partially supported by the Samsung Science and Technology Foundation (SSTF-BA2102-02) and the NRF grant RS-2025-00542968.
2. Preliminaries on Floer Theories
The proof of Theorem 1.2 involves two major theories: involutive Heegaard Floer homology [HM17] and filtered instanton Floer homology [NST24]. In this section, we briefly recall the basic concepts, key properties, and essential notions of each theory. For the first two subsections, we review the involutive Heegaard Floer homology and its local equivalence class in terms of monotone graded roots. In the third, we consider analogue in the Seiberg–Witten Floer theory. In the last, we briefly recall the filtered instanton homology and the construction of the $r_{s}$-invariants.
2.1. Involutive Heegaard Floer homology and local equivalence group
In this subsection, we review the involutive Heegaard Floer theory and the notion of local equivalence. We will only consider the minus flavor of the Heegaard Floer chain complex $CF^{-}(Y)$ [OS04], which is a graded $\mathbb{F}[U]$-module, where $U$ is a formal variable of grading 2-2.
Ozsváth and Szabó [OS04] defined the Heegaard Floer homology $HF^{-}(Y)$ for a closed 3-manifold $Y$, and proved that it is a diffeomorphism invariant. Hendricks and Manolescu [HM17] defined a homotopy involution $\iota$ on a Heegaard Floer chain complex by composing the map which arises from exchanging $\alpha$-curves and $\beta$-curves in the Heegaard diagram with the canonical chain homotopy equivalence based on the naturality [JTZ21]. Then they introduced the involutive Heegaard Floer homology $HFI^{-}(Y)$ as the homology of the mapping cone of $1+\iota$ on $CF^{-}(Y)$. When $Y$ is a homology sphere, $HFI^{-}(Y)$ gives rise to two homology cobordism invariants $\underline{d\mkern-2.0mu}\mkern 2.0mu(Y)$ and $\bar{d}(Y)$.
Instead of computing $HFI^{-}(Y)$ directly, Hendricks, Manolescu, and Zemke [HMZ18] defined an equivalence relation of the involutive Heegaard Floer chain complexes of homology spheres. The equivalence class is a homology cobordism invariant. Moreover, the classes form an abelian group $\mathfrak{I}$, called the local equivalence group, together with the homomorphism $h:\Theta_{\mathbb{Z}}^{3}\rightarrow\mathfrak{I}$ as mentioned in Section 1.
Let $Y_{1}$ and $Y_{2}$ be two homology spheres. Recall from [HM17] that a cobordism $W$ from $Y_{1}$ to $Y_{2}$ with a self-conjugate spinc-structure $\mathfrak{t}$ induces the $\iota$-equivariant $\mathbb{F}[U]$-chain map $F_{W,\mathfrak{t}}:CF^{-}(Y_{1})\rightarrow CF^{-}(Y_{2})$. Moreover, we have:
Theorem 2.1.
If two homology spheres $Y_{1}$ and $Y_{2}$ are homology cobordant, then there exist two grading-preserving $\iota$-equivariant $\mathbb{F}[U]$-chain maps $F:CF^{-}(Y_{1})\rightarrow CF^{-}(Y_{2})$ and $G:CF^{-}(Y_{2})\rightarrow CF^{-}(Y_{1})$ where $F$ and $G$ induce isomorphisms on homology after localization of $U$.
The local equivalence group is modeled on the homology cobordism group in the sense of Theorem 2.1.
Definition 2.2.
Let $C$ be a free, $\mathbb{Z}$-graded and finitely generated chain complex over $\mathbb{F}[U]$ with a grading-preserving $\mathbb{F}[U]$-equivariant homotopy involution $\iota$. The pair $(C,\iota)$ is called an $\iota$-complex if the localization of its homology is isomorphic to $\mathbb{F}[U,U^{-1}]$, i.e.,
$U^{-1}H_{*}(C)\cong\mathbb{F}[U,U^{-1}].$
Note that the involutive Heegaard Floer chain complex $(CF^{-}(Y),\iota)$ for a homology sphere $Y$ is an $\iota$-complex.
Definition 2.3.
Let $(C_{1},\iota_{1})$ and $(C_{2},\iota_{2})$ be two $\iota$-complexes. If there are grading-preserving $\iota$-equivariant chain maps $F:C_{1}\rightarrow C_{2}$ and $G:C_{2}\rightarrow C_{1}$ which induce isomorphisms on the homology after localization of $U$, then two $\iota$-complexes are called locally equivalent.
It is proved in [HMZ18] that the set of $\iota$-complexes modulo local equivalence with the operation tensor product forms an abelian group $\mathfrak{I}$, called the local equivalence group. The identity element of $\mathfrak{I}$ is given by the trivial complex consisting of a single $\mathbb{F}[U]$-tower starting at grading zero, together with the trivial $\iota$. Inverse elements in $\mathfrak{I}$ are obtained by dualizing $\iota$-complexes. Moreover, Hendricks, Manolescu, and Zemke [HMZ18] also proved that there is an $\iota$-equivariant chain homotopy equivalence:
$(CF^{-}(Y_{1}\#Y_{2}),\iota)\simeq((CF^{-}(Y_{1})\otimes CF^{-}(Y_{2}))[-2],\iota_{1}\otimes\iota_{2}),$
where [2][-2] denotes the grading shift by 2-2. By Theorem 2.1, the local equivalence class of $Y$ is a homology cobordism invariant, and from the connected sum formula above we have the homomorphism defined as:
$h:\Theta_{\mathbb{Z}}^{3}\to\mathfrak{I},\ \ h(Y)=(CF^{-}(Y),\iota)[-2],$
where the $d$-invariant [OS03a], the $\underline{d\mkern-2.0mu}\mkern 2.0mu$- and $\bar{d}$-invariants [HM17], and the $\phi_{n}$-invariants for all $n\geq 1$ [DHST23] factor through $\mathfrak{I}$.
2.2. Monotone graded subroots for Seifert fibered spheres
In this subsection, we explain the roles of graded roots and their monotone graded subroots to describe local equivalence classes of Seifert fibered spheres. We follow the references [Ném05, DM19, DS19, KŞ22].
In [Ném05], Némethi introduced two combinatorial objects to compute Heegaard Floer homology of Seifert fibered spheres effectively. The first object is called a graded root $(R,\nu)$ equipped with a grading function $\nu$ on the vertex set of $R$, where $R$ is an upwards-opening tree with an infinite downwards stem defined from the negative definite plumbing graph for a Seifert fibered sphere. The second one is a graded $\mathbb{F}[U]$-module, called the lattice homology $\mathbb{H}^{-}(R)$. It is combinatorially defined from $(R,\nu)$, whose grading is induced by $\nu$ with the formal variable $U$ of grading $2$. The grading on $\mathbb{H}^{-}(R)$ is denoted by the same symbol $\nu$ by abuse of notation. See [Ném05, DM19] for precise definitions.
As observed in [Dai18, DM19], for a Seifert fibered sphere, there is an involution $J$ on the vertex set of $R$ that reflects a graded root along the central vertical axis. We call the triple $(R,\nu,J)$ a symmetric graded root, and an example is depicted in Figure 1. Moreover, there is an $\mathbb{F}[U]$–equivariant action $J_{*}$ induced by $J$ on $\mathbb{H}^{-}(R)$. See [Dai18, Section 2.1] for further discussion.
Figure 1. The graded root $R$ with the involution $J$ for the Brieskorn sphere $\Sigma(3,4,13)$. The leaves and angles of $R$ are labeled by $v_{1},\ldots v_{6}$ and $\alpha_{1},\ldots\alpha_{5}$, respectively.
By combining the results of Ozsváth and Szabó [OS03b], Némethi [Ném05], and Dai and Manolescu [DM19], we have the following graded $\mathbb{F}[U]$-module isomorphism:
$F:HF^{-}(Y)\xrightarrow{\cong}\mathbb{H}^{-}(R)[\sigma]\quad\text{and}\quad F\circ\iota_{*}=J_{*}\circ F,$
where $\sigma$ denotes a grading shift [^1] For a Seifert fibered sphere, this shift is determined by algebraic topology of the minimal negative definite plumbing bounded by $Y$. See [KŞ22, Section 2.3] and references therein for a precise definition. and $\iota_{*}$ is the induced involution by $\iota$ on $HF^{-}(Y)$.
Following [DM19], we next construct a free and finitely generated $\mathbb{F}[U]$-chain complex whose homology is the same as $\mathbb{H}^{-}(R)$. See [DM19, Lemma 4.3] for a proof of this fact.
Let $v_{1},v_{2},\ldots,v_{n}$ (resp. $\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1}$) be the leaves (resp. the upward-opening angles) of $R$, enumerated from left to right in lexicographic order. The grading of vertex $v_{i}$ and the grading of the vertex supporting angle $\alpha_{i}$ are denoted by $\nu(v_{i})$ and $\nu(\alpha_{i})$, respectively; see Figure 1.
Figure 2. The standard complex $C_{*}(R)$ with the involution $J$ that captures the lattice homology $\mathbb{H}^{-}(\Sigma(3,4,13))$. Here, the solid (resp. the dashed) lines represent the action of $U$ (resp. $\partial$).
Then the generators of $C_{*}(R)$ are given as follows. For each leaf $v_{i}$, we place a single generator (denoted by also $v_{i}$ by abuse of notation) in grading $\nu(v_{i})$, so that we introduce an entire tower of generators $\mathbb{F}[U]v_{i}$. For each angle $\alpha_{i}$, we similarly place a single generator (denoted by also $\alpha_{i}$ by abuse of notation) in grading $\nu(\alpha_{i})+1$. We next define our differential to be identically zero on $v_{i}$, and set
$\partial\alpha_{i}=U^{(\nu(v_{i})-\nu(\alpha_{i}))/2}v_{i}+U^{(\nu(v_{i+1})-\nu(\alpha_{i}))/2}v_{i+1}$
for $\alpha_{i}$. Finally, we extend it to the entire complex linearly and $\mathbb{F}[U]$-equivariantly.
There is an involution $J$ on $C_{*}(R)$ induced by $J$ on $R$, and it is given by sending $v_{i}$ to $v_{n-i+1}$, $\alpha_{i}$ to $\alpha_{n-i}$, and extending linearly and $\mathbb{F}[U]$-equivariantly. See Figure 2. One can check that the pair $(C_{*}(R),J)$ forms an $\iota$-complex. Moreover, $J$ also induces the involution $J_{*}$ on $\mathbb{H}^{-}(R)$. See [DM19, Section 4] and [DS19, Section 2.3] for more details.
In [DM19, Section 6], Dai and Manolescu introduced the notion of a monotone graded subroot of a symmetric graded root $R$, to simply determine the local equivalence class of a Seifert fibered sphere.
Let $(R,\nu,J)$ be a symmetric graded root, and let $\mathcal{G}_{R}$ be the image of the grading function $\nu:R\to\mathbb{Z}$. For a positive integer $n$, let $h_{1},\ldots,h_{n}$ (resp. $r_{1},\ldots,r_{n}$) be a sequence of decreasing (resp. increasing) even integers in $\mathcal{G}_{R}$ such that $h_{n}\geq r_{n}$. Then a monotone graded subroot
$M=M(h_{1},r_{1};\ldots;h_{n},r_{n})$
of $R$ equipped with the same involution $J$ is constructed in the following fashion:
  1. (1)
    Form the stem of $M$ by drawing a single infinite tower with the uppermost vertex in grading $r_{n}$,
  2. (2)
    If $h_{n}>r_{n}$, use leaves $v_{i}$ and $Jv_{i}$ of $R$ to introduce leaves $v_{i}$ and $Jv_{i}$ of $M$ in grading $h_{i}$ for each $1\leq i<n$,
  3. (3)
    Connect $v_{i}$ and $Jv_{i}$ to the stem by using two paths meeting the stem in grading $r_{i}$ for each $1\leq i<n$,
  4. (4)
    If $h_{n}=r_{n}$, then set $v_{n}=Jv_{n}$ at grading $r_{n}$ in the second step.
By construction, the standard complex $(C_{*}(M),J)$ of a monotone graded subroot $M$ is a subcomplex of $(C_{*}(R),J)$, and it also forms an $\iota$-complex itself, see [DM19, Section 6]. However, it is sufficient to describe the local equivalence class of a Seifert fibered sphere completely, as proven by Dai and Manolescu.
Theorem 2.4 (Theorems 4.5 and 6.1, [DM19]).
Let $Y$ be a Seifert fibered sphere with a symmetric graded root $R$. Let $M$ be a monotone graded subroot of $R$. Then the $\iota$-complexes $(CF^{-}(Y),\iota)$, $(C_{*}(R),J)$, and $(C_{*}(M),J)$ are locally equivalent in $\mathfrak{I}$.
We now describe the parameterization of a monotone graded subroot for a Seifert fibered sphere by using involutive correction terms $\underline{d}$ and $\overline{d}$. See Figure 3 for an example.
Theorem 2.5 (Section 8, [DM19]; Theorem 4.4, [DS19]).
Let $Y$ be a Seifert fibered sphere with a graded root $R$. Then there exists a monotone graded subroot $M=M(h_{1},r_{1};\ldots;h_{n},r_{n})$ of $R$ such that
$h_{1}=\overline{d}(Y)\quad\text{and}\quad r_{n}=\underline{d}(Y).$
Figure 3. For the Brieskorn sphere $\Sigma(3,4,13)$, the monotone graded subroot M=M(0,2)M=M(0,-2) with the involution $J$ is drawn in black, compare with Figure 1.
By following the convention in [DS19], from now on, we will refer to the standard complex $(C_{*}(M),J)$ as just the monotone graded subroot $M$ for simplicity.
Now we present a key theorem, provided by Dai and Stoffregen [DS19], to prove Theorem 1.2. It describes how a monotone graded subroot decomposes into simpler ones up to local equivalence.
Theorem 2.6 (Theorem 4.2, [DS19]).
For any monotone graded subroot $M=M(h_{1},r_{1};\ldots;h_{n},r_{n})$, we have the local equivalence:
$M=\left(\sum_{i=1}^{n}M(h_{i},r_{i})\right)-\left(\sum_{i=1}^{n-1}M(h_{i+1},r_{i})\right).$
Relying on the above results, Dai and Stoffregen further proved the following theorem. Recall that $B(0)=\Sigma(2,3,5)$, and $B(n)=\Sigma(2n+1,4n+1,4n+3)$ for $n\geq 1$.
Theorem 2.7 (Theorem 1.1, [DS19]).
The image $\mathfrak{I}_{SF}$ of the subgroup $\Theta_{SF}^{3}$ generated by the Seifert fibered spheres under the homomorphism $h$ is infinitely generated, i.e.,
$\mathfrak{I}_{SF}=\langle h(B(n))\rangle\cong\mathbb{Z}^{\infty}.$
2.3. Local equivalence in Seiberg–Witten Floer theory
Based on the work of Manolescu [Man14], Stoffregen [Sto20] previously defined an analogous notion of the local equivalence class for homology spheres in Seiberg–Witten Floer theory, and computed these classes for Seifert fibered spheres. In this subsection, we present the equivalence between two notions of the local equivalence class in Seiberg–Witten Floer theory and the one in involutive Heegaard Floer theory for each Seifert fibered sphere.
As described in [Sto20], the local equivalence class of the $\operatorname{Pin}(2)$-equivariant Seiberg–Witten Floer stable homotopy type $SWF(Y)$ and of its associated chain complex of its spectrum also form abelian groups $\mathfrak{LE}$ and $\mathfrak{CLE}$, respectively, together with the homomorphisms:
$\Theta_{\mathbb{Z}}^{3}\rightarrow\mathfrak{LE}\rightarrow\mathfrak{CLE},$
where
  • the $\kappa$-invariant [Man14] and $\kappa o_{i}$-invariants [Lin15] for $i=0,\ldots,7$ factor through $\mathfrak{LE}$;
  • the $\delta$-invariant [Frø10], the $\alpha$-, $\beta$-, and $\gamma$-invariants [Man16], and the $\underline{\delta}$- and $\overline{\delta}$-invariants [Sto17b] factor through $\mathfrak{CLE}$.
See [Man18, Section 3] for further details.
Recall that $SWF(Y)$ recovers the $\operatorname{Pin}(2)$-equivariant Seiberg–Witten Floer homology of YY. [^2] Here, the Seiberg–Witten Floer homology means the $\operatorname{Pin}(2)$-equivariant Borel homology $\widetilde{H}_{*}^{\operatorname{Pin(2)}}(SWF(Y))$ of the space corresponding to $SWF(Y)$. By [LM18, Pan25], it turned out that this homology is isomorphic to the $\operatorname{Pin(2)}$-monopole Floer homology $\widecheck{HS_{*}}(Y)$ defined by [Lin18], based on [KM07]. Motivated from the facts that $\mathbb{H}^{-}(R)$ is isomorphic to $HF^{-}(Y)$ [Ném05], and hence to the Seiberg–Witten Floer homology [LM18, KLT20], Dai, Sasahira, and Stoffregen constructed the $\operatorname{Pin}(2)$-equivariant spectrum $\mathcal{H}(R)$ of the graded root $R$ of $Y$, and proved the following:
Theorem 2.8 (Theorems 1.2 and 7.3, [DSS23]).
Let $Y$ be a Seifert fibered sphere with a graded root $R$. Then there is a $\operatorname{Pin}(2)$-equivariant homotopy equivalence:
$\mathcal{H}(R)\simeq SWF(Y).$
In particular, for two Seifert fibered spheres $Y_{1}$ and $Y_{2}$, the following are equivalent:
  1. (1)
    $h(Y_{1})=h(Y_{2})$ in $\mathfrak{I}$.
  2. (2)
    $SWF(Y_{1})]=[SWF(Y_{2})]$ in $\mathfrak{LE}$.
2.4. Filtered instanton Floer homology and $r_{s}$-invariants
In this subsection, we briefly recall instanton Floer homology and its filtered version by Nozaki, Sato, and Taniguchi [NST24]. We follow the conventions and notations in [NST24].
For a given homology sphere $Y$, the Chern–Simons functional $cs_{Y}$ is defined on the space $A(Y)$ of $SU(2)$-connections on the product principal $SU(2)$-bundle $P_{Y}=Y\times SU(2)$. [^3] We write == rather than $\cong$ to emphasize the choice of the trivialization of the product bundle $P_{Y}$. This makes the choice of the reducible critical point $\theta$, which will be defined later, canonical. Since $cs_{Y}$ is invariant under the action of the group $\operatorname{Map}_{0}(Y,SU(2))$ consisting of degree 0 maps in the gauge group, $cs_{Y}:A(Y)\rightarrow\mathbb{R}$ factors through the configuration space $\widetilde{B}(Y)=A(Y)/\operatorname{Map}_{0}(Y,SU(2))$. We consider the subset $\widetilde{B}^{*}(Y)$ which consists of the orbits of irreducible connections, namely the stabilizer subgroup is $\{\pm I\}$ where $I$ denotes the constant map to the identity in $SU(2)$.
We fix a Riemannian metric on $Y$ and an appropriate perturbation $\pi$ of $cs_{Y}$. Then we construct an infinite-dimensional analogue of Morse chain complex of $\widetilde{B}^{*}(Y)$ using the perturbed Chern–Simon functional $cs_{Y,\pi}$. The set $\widetilde{R}^{*}_{\pi}(Y)$ of the irreducible critical points of $cs_{Y,\pi}$ in $\widetilde{B}^{*}(Y)$, consists of flat connections, namely, those with curvature $F_{A}=0$. The integer-valued Floer index indind, which is an analogue of the ordinary Morse index, is also well-defined on $\widetilde{R}^{*}(Y)$. Then the instanton Floer chain complex is defined as:
$CI_{i}(Y)=\mathbb{Z}\{a\in\widetilde{R}_{\pi}^{*}(Y)\ |\ ind(a)=i\},$
together with the boundary map $\partial:CI_{i}(Y)\rightarrow CI_{i-1}(Y)$ given by:
$\partial(a)=\sum\limits_{\mathclap{\begin{subarray}{c}b\in\widetilde{R}^{*}_{\pi}(Y)\\ \mathrm{ind}(b)=i-1\end{subarray}}}\#(M^{Y}(a,b)/\mathbb{R})b.$
Here, $M^{Y}(a,b)$ denotes the $1$-dimensional moduli space of instantons over the cylinder $\mathbb{R}\times Y$ where its restriction on $Y\times t$ is asymptotically $a$ when $t$ goes to -$\infty$ and $b$ when $t$ goes to $\infty$. Note that $\mathbb{R}$ acts on $M^{Y}(a,b)$ by the translation.
The boundary map satisfies $\partial^{2}=0$, so we obtain the instanton Floer homology $I_{*}(Y)$, which does not depend on the choice of Riemannian metric and perturbation, and is a diffeomorphism invariant of $Y$. By taking the dual complex $CI^{i}(Y)$, one can also obtain the instanton Floer cohomology $I^{*}(Y)$.
Note that the chain complex only involves irreducible critical points. Let $\theta\in A(Y)$ be the product $SU(2)$-connection on $P_{Y}$ so that $\theta$ is a reducible critical point. Donaldson [Don02] introduced an obstruction cocycle $\theta_{Y}:CI_{1}(Y)\rightarrow\mathbb{Z}$ in the first instanton Floer cochain complex $CI^{1}(Y)$ to count the trajectories from irreducible solutions to $\theta$ as follows:
$\theta_{Y}([a])=\#(M^{Y}(a,\theta)/\mathbb{R}),$
where $M^{Y}(a,\theta)$ is similarly defined to $M^{Y}(a,b)$. See [NST24] for a precise definition. The cocycle $\theta_{Y}$ forms a well-defined cohomology class $\theta_{Y}]\in I^{1}(Y)$, which vanishes whenever $Y$ bounds a negative definite $4$-manifold.
On the other hand, Fintushel and Stern [FS90] introduced the filtered instanton Floer homology $I_{*}^{[r,r+1]}(Y)$ for any fixed real number $r$ whose filtrations are given by the Chern–Simons functional. We recall the recent work of Nozaki, Sato, and Taniguchi [NST24] which extends the work of Fintushel and Stern to the filtered instanton Floer homology $I_{*}^{[s,r]}(Y)$ for a more general interval $[s,r]$. We write the set of real numbers:
$\Lambda_{Y}=cs_{Y}(\widetilde{R}_{\pi}(Y)),\quad\Lambda_{Y}^{*}=cs_{Y}(\widetilde{R}^{*}_{\pi}(Y)),\quad\text{and}\quad\mathbb{R}_{Y}=\mathbb{R}-\Lambda_{Y},$
where $\widetilde{R}_{\pi}(Y)$ is the set of the orbits of all critical points of $cs_{Y,\pi}$, whether irreducible or reducible.
Definition 2.9.
Let $s\in[-\infty,0]$, $r\in\mathbb{R}_{Y}$ and $\lambda_{Y}=\min\{|a-b||a\neq b,a,b\in\Lambda_{Y}\}$. The filtered instanton Floer chain complex is defined as:
$CI_{i}^{[s,r]}(Y)=\begin{cases}\mathbb{Z}\{a\in\widetilde{R}_{\pi}^{*}(Y)\ |\ ind(a)=i,cs_{Y,\pi}(a)\in(s,r)\}&\text{if $s\in\mathbb{R}_{Y}$,}\\ \mathbb{Z}\{a\in\widetilde{R}_{\pi}^{*}(Y)\ |\ ind(a)=i,cs_{Y,\pi}(a)\in(s-\lambda_{Y}/2,r)\}&\text{if $s\in\Lambda_{Y}$,}\end{cases}$
together with the boundary map given by the restriction of $\partial$ defined on the ordinary $CI_{i}(Y)$.
Since $\partial^{2}=0$, we have the filtered instanton Floer homology $I_{*}^{[s,r]}(Y)$, which is a diffeomorphism invariant of $Y$. Technically, the chain complex can be defined for arbitrary intervals, however, we avoid the case when $r\in\Lambda_{Y}$ to make the canonically defined map $i_{[s,r]}^{[s^{\prime},r^{\prime}]}:CI_{i}^{[s,r]}(Y)\rightarrow CI_{i}^{[s^{\prime},r^{\prime}]}(Y)$ behave well. More precisely,
Lemma 2.10 (Lemma 2.9, [NST24]).
Suppose $s\leq s^{\prime}\leq 0\leq r\leq r^{\prime}$ where $r,r^{\prime}\in\mathbb{R}_{Y}$. If both $r,r^{\prime}]$ and $s,s^{\prime}]$ do not contain any value of $\Lambda_{Y}^{*}$, then the map $i_{[s,r]}^{[s^{\prime},r^{\prime}]}:CI_{i}^{[s,r]}(Y)\rightarrow CI_{i}^{[s^{\prime},r^{\prime}]}(Y)$ defined by
$i_{[s,r]}^{[s^{\prime},r^{\prime}]}(a)=\begin{cases}a&\text{if $a\in CI_{i}^{[s^{\prime},r^{\prime}]}(Y)$,}\\ 0&\text{otherwise,}\end{cases}$
is a chain homotopy equivalence.
By taking duals of $CI_{i}^{[s,r]}(Y)$ and of the chain maps, we can also consider the filtered instanton Floer cochain complex $CI^{i}_{[s,r]}(Y)$ and its cohomology $I^{i}_{[s,r]}(Y)$, and the dual maps on them.
Let $Y_{1}$ and $Y_{2}$ be two homology spheres. Now we consider an oriented cobordism $W$ from $Y_{1}$ to $Y_{2}$. As in other Floer theories, $W$ induces a chain map between the instanton Floer chain complexes. In particular, when $b_{1}(W)=b_{2}^{+}(W)=0$, the (co)chain maps preserve the Floer indices:
$CW_{i}^{[s,r]}$ $:CI_{i}^{[s,r]}(Y_{1})\rightarrow CI_{i}^{[s,r]}(Y_{2});$
$CW^{i}_{[s,r]}$ $:CI^{i}_{[s,r]}(Y_{2})\rightarrow CI^{i}_{[s,r]}(Y_{1}).$
Let $\theta_{Y}^{[s,r]}$ be the restriction of $\theta_{Y}$ to $CI_{1}^{[s,r]}(Y)$. In [NST24], Nozaki, Sato, and Taniguchi observed that the map $\theta_{Y}^{[s,r]}$ defines a cocycle in $CI^{1}_{[s,r]}(Y)$. The cohomology class $\theta_{Y}^{[s,r]}]\in I^{1}_{[s,r]}(Y)$ also behaves well under the induced map $IW^{1}_{[s,r]}$ by $CW^{1}_{[s,r]}$ on the first cohomology. More precisely,
Theorem 2.11 (Lemma 2.13, [NST24]).
If there is a negative definite cobordism $W$ from $Y_{1}$ to $Y_{2}$ with $H^{1}(W;\mathbb{R})=0$, then for $r\in\mathbb{R}_{Y_{1}}\cap\mathbb{R}_{Y_{2}}$,
$IW^{1}_{[s,r]}([\theta_{Y_{2}}^{[s,r]}])=|H_{1}(W;\mathbb{Z})|[\theta_{Y_{1}}^{[s,r]}].$
In particular, if $Y$ bounds a negative definite $4$-manifold, then the obstruction class $\theta_{Y}^{[s,r]}]\in I^{1}_{[s,r]}(Y)$ vanishes.
Let $i_{[s^{\prime},r^{\prime}]}^{[s,r]}$ be the dual map of $i^{[s^{\prime},r^{\prime}]}_{[s,r]}$. We use the same notation for the induced maps on (co)homology by abusing notation. One can observe from Lemma 2.10 that:
Lemma 2.12 (Lemma 2.15, [NST24]).
Suppose $s\leq s^{\prime}\leq 0\leq r\leq r^{\prime}$ where $r,r^{\prime}\in\mathbb{R}_{Y}$. Then,
$i_{[s^{\prime},r^{\prime}]}^{[s,r]}[\theta_{Y}^{[s^{\prime},r^{\prime}]}]=[\theta_{Y}^{[s,r]}].$
[…]
Definition 2.13 (Definition 3.2, [NST24]).
Let $s\in[-\infty,0]$. The $r_{s}$-invariant of $Y$ is defined as:
$r_{s}(Y)=\sup\{r\in(0,\infty]\ |\ [\theta_{Y}^{[s,r]}\otimes id_{\mathbb{Q}}]=0\in I^{1}_{[s,r]}(Y;\mathbb{Q})\}.$
By definition and the vanishing property of the obstruction class, if $Y$ bounds a negative definite $4$-manifold, then $r_{s}(Y)=\infty$ for all $s\in[-\infty,0]$. In particular, for any $s\in[-\infty,0]$ and any Seifert fibered sphere $\Sigma(a_{1},\ldots,a_{n})$, we have
$r_{s}(\Sigma(a_{1},\ldots,a_{n}))=\infty.$
Nozaki, Sato, and Taniguchi [NST24, Theorem 5.15] proved that the value $r_{s}(Y)$ is invariant under rational homology cobordism. They also obtained the following connected sum formula [NST24, Theorem 1.1(4)]
$r_{s_{1}+s_{2}}(Y_{1}\#Y_{2})\geq\min\{r_{s_{1}}(Y_{1})+s_{2},\ r_{s_{2}}(Y_{2})+s_{1}\},$
so it is useful to take $s=0$ to obtain linear independence in $\Theta_{\mathbb{Q}}^{3}$.
Theorem 2.14 (Corollary 5.6, Theorem 5.15, [NST24]).
Let $\{Y_{i}\}_{i\geq 1}$ be an infinite family of homology spheres. If all $r_{0}(Y_{i})$ are distinct and finite, and $r_{0}(-Y_{i})=\infty$, then $\{Y_{i}\}_{i\geq 1}$ is linearly independent in $\Theta_{\mathbb{Q}}^{3}$.
Theorem 2.14 will play a crucial role in the proof of Theorem 1.2, coupled with the useful formula for $r_{0}(Y)$ for a Seifert fibered sphere $Y$ such that $R(Y)>0$, where $R$ is the integer defined for a Seifert fibered sphere by Fintushel and Stern [FS85]. We close this section by presenting such a formula.
Theorem 2.15 (Corollary 1.4, [NST24]).
Let $Y$ be the Seifert fibered sphere $\Sigma(a_{1},\ldots,a_{n})$. If $R(Y)>0$, then for any $s\in[-\infty,0]$,
$r_{s}(-Y)=\frac{1}{4a_{1}\cdots a_{n}}\cdot$
3. Proof of Theorem 1.2
In this section, we prove Theorem 1.2. To do so, we first present explicit linear dependence relations of $h(Y_{1}(n))$, $h(Y_{2}(n))$ and $h(Y_{3}(n))$ in terms of $h(B(n))$ in the local equivalence group $\mathfrak{I}$ using Theorem 2.6. Then we prove linear independence of them in $\Theta_{\mathbb{Q}}^{3}$ by using Theorems 2.14 and 2.15.
It is well-known that the graded root of the Poincaré homology sphere has the monotone graded subroot $M(2,2)$ since $\overline{d}(\Sigma(2,3,5))=2=\underline{d}(\Sigma(2,3,5))$. By Theorem 2.5, $M(2,2)$ is a single tower with the uppermost vertex in grading $2$.
We now describe the monotone graded subroots of $B(n)$ for $n\geq 1$ [DS19] (see also [Sto17a]) and the ones for the following infinite families [KŞ22] (see also [KŞ20]):
  • $Y_{1}(n)=\Sigma(4n+1,6n+2,12n+1)$,
  • $Y_{2}(n)=\Sigma(4n-1,6n-2,12n-1)$,
  • $Y_{3}(n)=\Sigma(8n+1,12n+1,24n+5)$.
Theorem 3.1 (Theorem 1.1, [DS19]; Theorem 1.2, [KŞ22]).
For $n\geq 1$, the Brieskorn spheres $B(n)$, $Y_{1}(n)$ and $Y_{2}(n)$ have the monotone graded subroots
$M(2n,0),\quad M(4n,0;2n,2n),\quad\text{and}\quad M(4n-2,0;2n,2n),$
respectively, which are depicted in Figure 4. Moreover, the Brieskorn spheres $Y_{3}(n)$ has the same monotone graded subroot as $B(4n)$.
Figure 4. The monotone graded subroots of $B(n)$, $Y_{1}(n)$ and $Y_{2}(n)$ for $n\geq 1$.
Lemma 3.2.
In the local equivalence group $\mathfrak{I}$, we have the following for $n\geq 1:$
  1. (1)
    $h(Y_{1}(n))=h(B(2n)\ \#\ -B(n)\ \#\ nB(0))$,
  2. (2)
    $h(Y_{2}(n))=h(B(2n-1)\ \#\ -B(n)\ \#\ nB(0))$,
  3. (3)
    $h(Y_{3}(n))=h(B(4n))$.
Proof.
By applying Theorems 2.4 and 3.1 together, we first see that
$h(Y_{1}(n))=M(4n,0;2n,2n)\quad\text{and}\quad h(Y_{2}(n))=M(4n-2,0;2n,2n).$
Note that $M(2n,2n)$ is the single tower with the uppermost vertex in grading $2n$, and it is obtained by tensoring $n$ copies of $M(2,2)$, where $M(2,2)=h(B(0))$. So we have:
$M(2n,2n)=\underbrace{M(2,2)+\cdots+M(2,2)}_{n\text{ copies}}.$
We next use Theorem 2.6 to decompose the monotone graded subroots above. Since $h:\Theta^{3}_{\mathbb{Z}}\to\mathfrak{I}$ is a homomorphism, for the case (1), we have:
$h(Y_{1}(n))$ $=M(4n,0;2n,2n),$
$=M(4n,0)+M(2n,2n)-M(2n,0),$
$=M(4n,0)+\underbrace{M(2,2)+\cdots+M(2,2)}_{n\text{ copies}}-M(2n,0),$
$=h(B(2n)\ \#\ -B(n)\ \#\ nB(0)).$
One can similarly prove the case (2). Finally, the case (3) is a direct consequence of Theorem 3.1. ∎
Next, we will show that all the Brieskorn spheres $B(n)$, $Y_{1}(n)$, and $Y_{2}(n)$ are linearly independent up to rational homology cobordism. In particular, the homology spheres in Theorem 1.2 are generated by the above.
Lemma 3.3.
The homology spheres below are linearly independent in $\Theta^{3}_{\mathbb{Q}}:$
$\big{\{}B(n)\big{\}}_{n\geq 0}\ \cup\ \big{\{}Y_{1}(n)\big{\}}_{n\geq 1}\ \cup\ \big{\{}Y_{2}(n)\big{\}}_{n\geq 1}.$
Proof.
Since every Brieskorn sphere bounds a negative definite $4$-manifold, by Theorem 2.11 and definition of the $r_{s}$-invariants,
$r_{0}(B(n))=r_{0}(Y_{1}(n))=r_{0}(Y_{2}(n))=\infty.$
By [KŞ20, Proposition 4.3], all of these homology spheres have negative definite plumbing graphs with central weights e=2e=-2. Thus, since R(Y)=2e3R(Y)=-2e-3 for any Seifert fibered sphere $Y$ [NZ85], we have:
$R(B(n))=R(Y_{1}(n))=R(Y_{2}(n))=1.$
By Theorem 2.15, $r_{0}(-\Sigma(p,q,r))=\frac{1}{4pqr}$, provided $R(\Sigma(p,q,r))>0$. Thus, according to Theorem 2.14, it suffices to show that all the families of Brieskorn spheres have distinct product of the exponents. It is clear that $(2n+1)(4n+1)(4n+3)$, $(4n+1)(6n+2)(12n+1)$ and (4n1)(6n2)(12n1)(4n-1)(6n-2)(12n-1) are never the same as $2\cdot 3\cdot 5=30$, which implies that
$r_{0}(-B(0))$ $\neq r_{0}(-Y_{1}(m)),$
$r_{0}(-B(0))$ $\neq r_{0}(-Y_{2}(m)).$
For the family $B(n)$ with $n>0$, note that $(2n+1)(4n+1)(4n+3)$ is always odd. In contrast, we see that $(4m\pm 1)(6m\pm 2)(12m\pm 1)$ are always even. Thus, there are no pair of positive integers $(n,m)$ such that $(2n+1)(4n+1)(4n+3)=(4n\pm 1)(6m\pm 2)(12m\pm 1)$. Thus,
$r_{0}(-B(n))$ $\neq r_{0}(-Y_{1}(m)),$
$r_{0}(-B(n))$ $\neq r_{0}(-Y_{2}(m)).$
Finally, by dividing both of $(4m\pm 1)(6m\pm 2)(12m\pm 1)$ by $2$, define $f(a)=(4a+1)(3a+1)(12a+1)$ and g(b)=(4b1)(3b1)(12b1)g(b)=(4b-1)(3b-1)(12b-1). Then for any $a\geq 1$, we can see that $g(a)<f(a)<g(a+1)$. Since $g(b)$ is strictly increasing when $b\geq 1$, there cannot exist a pair of positive integers $(a,b)$ such that $f(a)=g(b)$, which implies that
$r_{0}(-Y_{1}(m))\neq r_{0}(-Y_{2}(k)).$
Now since $r_{0}(B(n))=r_{0}(Y_{1}(n))=r_{0}(Y_{2}(n))=\infty$, and $r_{0}(-B(n))$, $r_{0}(-Y_{1}(m))$ and $r_{0}(-Y_{2}(k))$ are finite and distinct, all of them are linearly independent in $\Theta_{\mathbb{Q}}^{3}$ by Theorem 2.14. ∎
For the same reason as in the proof of Lemma  3.3, we see that $R(Y_{3}(n))=1$. Since $(8n+1)(16n+1)(16n+3)\not\equiv(8n+1)(12n+1)(24n+5)$ mod $4$, we have that
$r_{0}(-B(4n))\neq r_{0}(-Y_{3}(n)).$
Thus, we have the following lemma as a consequence of Theorems 2.14 and 2.15:
Lemma 3.4.
The homology spheres below are linearly independent in $\Theta_{\mathbb{Q}}^{3}:$
$\big{\{}B(4n)\big{\}}_{n\geq 0}\ \cup\ \big{\{}Y_{3}(n)\big{\}}_{n\geq 1}.$
Now we are ready to prove Theorem 1.2.
Proof of Theorem 1.2.
Consider the homology spheres of the forms
$\big{\{}Y_{1}(n)\ \#\ -B(2n)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}\cup\big{\{}Y_{2}(n)\ \#\ -B(2n-1)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}\cdot$
By (1) and (2) in Lemma 3.2, we know that they lie in the kernel of the homomorphism $h:\Theta^{3}_{\mathbb{Z}}\to\mathfrak{I}$. Applying Lemma 3.3, we also see that they are linearly independent in $\Theta^{3}_{\mathbb{Q}}$, which completes the proof.
Consider the other family of homology spheres
$\{Y_{3}(n)\#-B(4n)\}_{n\geq 1}.$
[…]